# When Can Model-Free Reinforcement Learning be Enough for Thinking?

### v1

@@ Abstract
Recent work on large language models has demonstrated the use of model-free reinforcement learning (RL) to train reasoning-like capabilities. The emergence of "thinking" through model-free RL is interesting as thinking actions neither produce reward nor change the external world state to one where the agent is more likely to get reward. This paper seeks to build a domain-independent understanding of when model-free RL will lead to "thinking" as a strategy for reward maximization. To build this understanding, we first introduce a theoretical model which we call a thought Markov decision process (MDP). Thought MDPs minimally extend the classical MDP model to include an abstract notion of thought state and thought action. Using the thought MDP model, we prove the importance of policy initialization in determining whether or not thinking emerges and show formally that thought actions are equivalent to the agent choosing to perform a step of policy improvement before continuing to act. We then show that open-source LLMs satisfy the conditions that our theory predicts are necessary for model-free RL to produce thinking-like behavior. Finally, we hypothesize sufficient conditions that would enable thinking to be learned outside of language generation and introduce a toy domain where a combination of multi-task pre-training and designated thought actions enable more data-efficient RL compared to non-thinking agents.
@@ 1 Introduction
Informally, we define thinking to be actions that do not directly produce reward or affect the external state of an agent’s environment but that lead the agent to take a course of action that increases the reward it will receive in the future. To answer our research question, we first formulate a minimal extension of the classical MDP model to explicitly model thought actions as well as a notion of a controllable thought state. We then show how policy initialization plays a central role in whether policy iteration will select thought actions. Under our theoretical model, we will show that thinking can be viewed as selecting between a set of sub-policies that are already contained in the learning agent’s policy function and that thought actions can be interpreted as the agent choosing to run one or more steps of policy improvement before continuing to act. We then discuss how LLMs instantiate our thought MDP formalism and provide empirical evidence that they exhibit the necessary conditions for thinking to emerge. Our final contribution is to introduce a simple domain and multi-task pre-training set-up that induces the condition under which model-free RL will discover thinking behavior. This simple domain provides a basis for future work studying agents that learn to think and act. We conclude by discussing open questions and directions for future work raised by this model of deliberative thinking in RL.
@@ 2 Related Work
A hallmark of System 2 processing is deliberative planning as opposed to reflexive action. Thus, a natural way to instantiate System 2 in artificial agents is to use decision-time planning, e.g., Monte Carlo tree search (Kocsis and Szepesvári, 2006), and many works in the RL literature have used this approach (e.g., Anthony et al., 2017; Silver et al., 2016, 2017; Silver, 2009; Schrittwieser et al., 2021; Shah et al., 2020). Decision-time planning is particularly effective when the agent’s policy or value function is sub-optimal but the agent possesses an accurate model of state transitions (Sutton and Barto, 2018). In this case, planning can be viewed as focused policy improvement to improve the decision at the agent’s current state before it acts. As we will show, the idea of thinking as local policy improvement also applies when an RL agent learns to think in a thought MDP. The key difference is that we consider a more abstract model of thinking without any notion of forward prediction and search. There is also no natural answer to the question of when to stop planning and act with most decision-time planning methods. MCTS is an anytime algorithm and the general practice is to allow as much search is possible in the application domain (Silver et al., 2016, 2017). On the other hand, model-free learning in our thought MDP model will directly relate the decision about when to stop thinking to the objective of return maximization. Some work in the (non-learning) planning literature has also studied the question of when to think vs. when to act (Cashmore et al., 2019; Coles et al., 2024).
@@
As neural networks are general-purpose function approximators, they can, in principle, learn algorithmic planning or reasoning computations and prior work has attempted to induce planning capabilities through specialized architectures (Tamar et al., 2017; Farquhar et al., 2018; Guez et al., 2019, 2018; Niu et al., 2018; Weber et al., 2018; Sykora et al., 2020; Schleich et al., 2019; Oh et al., 2017). Guez et al. showed that planning-like computations could emerge through model-free learning (2019). Bush et al. (2025) applied mechanistic interpretability approaches to show that the approach of Guez et al. (2019) did in fact learn planning computations (i.e., the network internally learned to propose and evaluate future candidate action sequences). While such works also show that model-free RL can produce thinking behavior, the main difference is that thinking refers to the computation done by the agent’s policy, whereas we study thinking as an action selected by the agent’s policy. More similar to this conception of thinking, Chung et al. (2023) introduce the thinker architecture in which a policy is trained to select actions both for environment interaction and planning. However, they conduct thinking for a fixed number of steps per environment step and thus the agent is learning how to think rather than when to think.
@@
Other Forms of Mental Action
The idea of thinking as a form of action has a long history in AI and RL. Minsky theorized on the mind as a society of agents whose collective actions produce cognition and action (Minsky, 1986). Klopf’s hedonistic neuron hypothesis modelled neurons as individual RL agents (Klopf, 1982). This hypothesis has led to a line of work studying alternative neural network architectures in which artificial neural activity is produced by the stochastic policies of simple RL agents (Thomas, 2011; Kostas et al., 2019; Gupta et al., 2021). More similar to our work, some prior work has considered augmenting the agent’s environment action-space with a form of mental action. These works have focused on the challenge of memory in partially observable domains and using explicit memory read or write actions as an alternative to recurrent neural networks. For instance, Peshkin et al. (2001) augment an RL agent with a set of memory write actions and the contents of the memory are then provided to the agent as an additional input along with the environment state; Zhang et al. (2015) extend this approach to robot manipulation tasks. Various neural architectures have been developed with external memory that can be written to and read from (e.g., Graves et al., 2014; Zaremba and Sutskever, 2016) and Oh et al. (2016) explored these architectures for RL. The thought MDP model differs in that the agent has access to a Markov state representation and so thought actions serve to manipulate the agent’s policy as opposed to remembering details of the past.
@@ 3.1 RL in Markov Decision Processes
RL environments are typically modeled as Markov decision processes (MDPs). Formally, an MDP is a tuple, $\langle\mathcal{S},\mathcal{A},p,r,\gamma\rangle$, where $\mathcal{S}$ is the environment state space, $\mathcal{A}$ is the set of actions that the agent can take to influence the environment state, $p:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is a stochastic state transition function, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a scalar-valued reward function, and $\gamma$ is a discount factor. At any moment in time, the agent is in state $s_{t}$, takes an action, $a_{t}$, receives a reward, and transitions to a new state. Then the process repeats from $s_{t+1}$ until a terminal state, $s_{\infty}$, is reached. The agent selects actions according to a policy, $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$. The value of using a policy from a particular state, $s$, is defined to be $v_{\pi}(s)\coloneqq\mathbf{E}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{t}\sim\pi(\cdot|s_{t})]$. In RL, the agent’s objective is to find a policy that maximizes $v_{\pi}(s)$ in all states.
@@ 3.2 Thought MDPs
We now extend the MDP model to explicitly model thinking. We formally define a thought MDP as the tuple $\langle\mathcal{S},\mathcal{A},p,r,\gamma,\mathcal{T},\mathcal{C},p_{\mathcal{T}}\rangle$, where $\mathcal{S},\mathcal{A},p,r,\gamma$ are defined as they are for MDPs. We add $\mathcal{T}$ as the set of thought states, $\mathcal{C}$ as the set of thought actions, and $p_{\mathcal{T}}:\mathcal{S}\times\mathcal{T}\times\mathcal{C}\rightarrow\Delta(\mathcal{T})$ as the thought transition function. Recalling our informal definition of thinking in the introduction, we emphasize that thought states and actions do not affect environment state transitions and rewards. The agent’s objective remains to maximize cumulative discounted reward across all environment states.
There are different ways to define the agent’s policy in a thought MDP. In this work, we formalize policies as a mapping $\pi:\mathcal{S}\times\mathcal{T}\rightarrow\Delta(\mathcal{A}\cup\mathcal{C})$. This choice means that the agent can select either an environment action or a thought action at any interaction time-step but not both. We make this choice to better connect to work in LLMs, but an alternative is that the policy is a mapping $\pi:\mathcal{S}\times\mathcal{T}\rightarrow\Delta(\mathcal{A}\times\mathcal{C})$, i.e., the agent can think and act at the same time. Such modelling might be more useful in real-time domains where the agent cannot simply sit and think as the rest of the world evolves around it. There, of course, may be other alternatives that could be considered in future work. Below, we will use the notation $\pi(\tau)$ to refer to the agent’s state-dependent policy with the thought state fixed at $\tau$.
In a thought MDP, interaction proceeds as follows. Episodes begin in an environment state, $s_{0}$, and thought state, $\tau_{0}$, and the agent chooses either an environment action or thought action according to its policy. If $a_{0}\in\mathcal{A}$ then the environment state, $s_{1}$, at the next time-step is sampled from $p(s_{0},a_{0})$, the thought state $\tau_{1}$ keeps the value of $\tau_{0}$, and the agent receives reward $r_{0}\coloneqq r(s_{0},a_{0})$. Conversely, if $a_{0}\in\mathcal{C}$, then $s_{1}$ keeps the value of $s_{0}$, $\tau_{1}$ is sampled from $p_{\mathcal{T}}(s_{0},\tau_{0},a_{0})$, and the agent receives $r_{0}=0$. Then the process repeats until the agent reaches $s_{\infty}$, with the agent selecting either an environment or thought action from its policy. For thought MDPs, we define the value function to be $v_{\pi}(s,\tau)\coloneqq\mathbf{E}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,\tau_{0}=\tau,a_{t}\sim\pi(\cdot|s_{t},\tau_{t})].$
@@
Our formalism raises important questions about time. First, should the discount factor be applied equally for both thinking and non-thinking time-steps? Equal application discourages thinking as thought actions do not influence reward either directly or indirectly through the environment state. Nevertheless, as we shall see, thought actions still might be selected if they ultimately cause the agent to choose a better environment action. Alternatively, we could apply a different discount factor for thinking time-steps to reflect the actual time-delay of thinking compared to acting in a given domain. For example, if thinking takes $(1/k)$ the time of any environment action then we could use a discount of $\gamma^{\frac{1}{k}}$. In this work, we will assume that the discount is applied the same at both thinking and non-thinking time steps. The second related issue is that the proposed model assumes that the environment state remains constant while the agent takes thought actions. Such an assumption might be reasonable for relatively static environments (such as generating text or board games) but is problematic in dynamic environments (such as autonomous driving) where waiting to act can be consequential. Again, for simplicity, we will focus on the static environment case, but note this issue as another interesting direction to refine the model.
@@
The proof is by contradiction. Suppose $\pi^{\star}$ is an optimal policy and $\exists s\in\mathcal{S},\tau\in\mathcal{T}$ such that $\pi(s,\tau)\in\mathcal{C}$. Because thought actions cannot produce reward, the optimal policy must eventually reach a thought state, $\tilde{\tau}$, where $\pi^{\star}(s,\tilde{\tau})\in\mathcal{A}$. If that happens after $k$ thought actions, then $v_{\pi^{\star}}(s,\tau)=\gamma^{k}v_{\pi^{\star}}(s,\tilde{\tau})$. But then the policy could be strictly improved by setting $\pi^{\star}(s,\tau)\leftarrow\pi^{\star}(s,\tilde{\tau})$. This contradicts the assumption that $\pi^{\star}$ was optimal and completes the proof. ∎
@@ 4 Policy Initialization Determines Emergence of Thinking
For exposition’s sake, suppose $\mathcal{T}=\{\tau_{0},\tau_{1}\}$, in which case we can view the agent’s policy, $\pi$, as consisting of two sub-policies, $\pi(\tau_{0})$ and $\pi(\tau_{1})$. Now consider the example depicted in Figure 1 where $\pi(\tau_{0})$ is initialized sub-optimally (always takes an action that leads away from the goal) while $\pi(\tau_{1})$ is initialized to be optimal. We show the progress of policy iteration in this domain in Figure 1. After the first iteration of policy improvement, the policy has learned to first change the agent’s thought state to $\tau_{1}$ (top row) from which it then follows $\pi(\tau_{1})$ to reach the goal. After four iterations, in environment states close to the goal, the policy just directly moves right without changing its thought state while continuing to first take a thought action in states that are far from the goal. Finally, after ten iterations, the policy converges to the optimal policy and simply moves to the right without thinking. This example shows that, while thinking is sub-optimal in the long-run, it can be beneficial early in learning by allowing the agent to use sub-policies already contained in its policy function.
@@
Policy iteration sets $\pi^{\prime}(s,\tau)\leftarrow\arg\max_{x\in\mathcal{A}\cup\mathcal{C}}q_{\pi}(s,\tau,x)$ where
$q_{\pi}(s,\tau,x)\coloneqq\begin{cases}r(s,x)+\gamma\mathbf{E}_{s^{\prime}}[v_{\pi}(s^{\prime},\tau)]&\text{if }x\in\mathcal{A}\\ \gamma v_{\pi}(s,\tau^{\prime})&\text{if }x\in\mathcal{C}.\end{cases}$
@@
Thus, a thought action is selected when the policy is such that $\pi(s,\tau^{\prime})$ is expected to lead to more reward than $\pi(s,\tau)$. ∎
Although, Theorem 5 does not directly say anything about policy initialization, the condition for a thought action to be selected requires $\pi(s,\tau)$ and $\pi(s,\tau^{\prime})$ to be somehow set up as different such that there could be an advantage in shifting from $\tau$ to $\tau^{\prime}$. Thus, policy initialization is critical for whether thinking will emerge or not.
Theorem 5 deals with exact policy iteration and does not address the impact of policy initialization on the emergence of thinking when learning from samples. We next turn to the concept of the effective horizon introduced by Laidlaw et al. (2023). The effective horizon was introduced as a problem complexity parameter for MDPs that aligns with the observed difficulty of common RL benchmarks. Laidlaw et al. then derived sample complexity bounds for model-free RL algorithms in terms of the effective horizon and showed that these bounds were predictive of the success of deep RL algorithms on benchmark problems. We will build upon their specific result on goal MDPs.
@@
A goal (thought) MDP is an (thought) MDP with a set of absorbing states, $\mathcal{S}_{\mathtt{goal}}$, and $r(s,a)=1$ if $s\in\mathcal{S}_{\mathtt{goal}}$ and 00 otherwise.
For goal MDPs, Laidlaw et al. (2023) showed that the effective horizon, $H$ can be upper-bounded by $1+\log_{\mathcal{A}}\frac{\log 2T}{p}$ where $p\leq\Pr_{\mathtt{expl}}(s_{T}\in\mathcal{S}_{\mathtt{goal}}|s_{t},a_{t})$ is a lower bound on the probability of the initial exploration policy, $\pi_{\mathtt{expl}}$, discovering a goal state after taking action $a$ in state $s$ at any time-step. Intuitively, if $p$ is larger, then an RL algorithm has an easier time discovering rewarding action sequences. We next show that thought actions can be used to reduce the effective horizon of a given instance. Note that Laidlaw et al. (2023) proved their results assuming deterministic MDPs, so we also adopt this assumption for the next proposition.
@@
Suppose we have a goal thought MDP that has $\mathcal{T}=\{\tau_{0},\tau_{1}\}$. Let the initial policy, $\pi$, be such that there is some $p_{0}$ such that $\Pr(s_{T}\in s_{\mathtt{goal}}|s,a,\pi(\tau_{0}))>p_{0}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ and some $p_{1}$ such that $\Pr(s_{T}\in s_{\mathtt{goal}}|s,a,\pi(\tau_{1}))>p_{1}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. Let $c_{s}$ be the thought action such that $\tau_{1}=p_{\mathcal{T}}(s,\tau_{0},c_{s})$ and let $p_{c}$ lower bound the probability $\pi(c_{s}|s,\tau_{0})$. If $p_{c}\cdot p_{1}>p_{0}$, then thought actions reduce the upper bound on the effective horizon of the MDP.
See Appendix A for the proof. Intuitively, policy $\pi(\tau_{1})$ has a higher probability of finding a goal state from any state compared to simply running $\pi(\tau_{0})$. If the probability of taking a thought action that changes $\tau_{0}$ to $\tau_{1}$ is not too low then it is easier to find a goal state by first changing the thought state and then executing $\pi(\tau_{1})$. Again, this result shows that the benefit of thinking depends crucially on policy initialization.
@@
One interpretation of Theorem 5 is that thought actions can function as policy improvement operators applied to a particular state. This interpretation is interesting as it aligns with the use of decision-time planning in RL to refine an agent’s choice of action in a way that focuses computation on its current state (Sutton and Barto, 2018). While thought actions do not involve look-ahead search, Theorem 5 shows that their utility is also in providing local policy improvement. Theorem 5 shows this utility for the choice of a single thought action and we also present a corollary showing that if policy improvement produces a policy that thinks for consecutive steps in $s$ then each thinking step will further improve upon the action $\pi(s,\tau)$.

### v2

@@ Abstract
Recent work on large language models has demonstrated the use of model-free reinforcement learning (RL) to train reasoning-like capabilities. The emergence of "thinking" through model-free RL is interesting as thinking actions neither produce reward nor change the external world state to one where the agent is more likely to get reward. This paper seeks to build a domain-independent understanding of when model-free RL will lead to such "thinking" as a strategy for reward maximization. To build this understanding, we first introduce a theoretical model which we call a thought Markov decision process (MDP). Thought MDPs minimally extend the classical MDP model to include an abstract notion of thought state and thought action. Using the thought MDP model, we prove the importance of policy initialization in determining whether or not thinking emerges and show formally that thought actions are equivalent to the agent choosing to perform a step of policy improvement before continuing to act. We then show that open-source LLMs satisfy the conditions that our theory predicts are necessary for model-free RL to produce thinking-like behavior. Finally, we hypothesize sufficient conditions that would enable thinking to be learned outside of language generation and introduce a toy domain where a combination of multi-task pre-training and designated thought actions enable more data-efficient RL compared to non-thinking agents.
@@ 1 Introduction
Informally, we define thinking to be actions that do not directly produce reward or affect the external state of an agent’s environment but that lead the agent to take a course of action that increases the reward it will receive in the future. To answer our research question, we first formulate a minimal extension of the classical MDP model to explicitly model thought actions and a notion of a controllable thought state. We then show how policy initialization plays a central role in whether policy iteration will select thought actions. Under our theoretical model, we will show that thinking can be viewed as selecting between a set of sub-policies that are already contained in the learning agent’s policy function and that thought actions can be interpreted as the agent choosing to run one or more steps of policy improvement before continuing to act. We then discuss how LLMs instantiate our thought MDP formalism and provide empirical evidence that they exhibit the necessary conditions for thinking to emerge. Our final contribution is to introduce a simple domain and multi-task pre-training set-up that induces the condition under which model-free RL will discover thinking behavior. This simple domain provides a basis for future work studying agents that learn to think and act. We conclude by discussing open questions and directions for future work raised by this model of deliberative thinking in RL.
@@ 2 Related Work
A hallmark of System 2 processing is deliberative planning as opposed to reflexive action. Thus, a natural way to instantiate System 2 in artificial agents is to use decision-time planning, e.g., Monte Carlo tree search (Kocsis and Szepesvári, 2006), and many works in the RL literature have used this approach (e.g., Anthony et al., 2017; Silver et al., 2016, 2017; Silver, 2009; Schrittwieser et al., 2021; Shah et al., 2020). Decision-time planning is particularly effective when the agent’s policy or value function is sub-optimal but the agent possesses an accurate model of state transitions (Sutton and Barto, 2018). In this case, planning can be viewed as focused policy improvement to improve the decision at the agent’s current state before it acts. As we will show, the idea of thinking as local policy improvement also applies when an RL agent learns to think in our new thought MDP model. The key difference is that we consider a more abstract model of thinking without any notion of forward prediction and search. Furthermore, there is also no natural answer to the question of when to stop planning and act with most decision-time planning methods. MCTS is an anytime algorithm and the general practice is to allow as much search is possible in the application domain (Silver et al., 2016, 2017). On the other hand, model-free learning in our thought MDP model will directly relate the decision about when to stop thinking to the objective of return maximization. Some work in the (non-learning) planning literature has also studied the question of when to think vs. when to act (Cashmore et al., 2019; Coles et al., 2024).
@@
As neural networks are general-purpose function approximators, they can, in principle, learn algorithmic planning or reasoning computations and prior work has attempted to induce planning capabilities through specialized architectures (Tamar et al., 2017; Farquhar et al., 2018; Guez et al., 2019, 2018; Niu et al., 2018; Weber et al., 2018; Sykora et al., 2020; Schleich et al., 2019; Oh et al., 2017). Guez et al. showed that planning-like computations could emerge through model-free learning (2019) and Bush et al. (2025) applied mechanistic interpretability approaches to show that the approach of Guez et al. (2019) did in fact learn planning computations (i.e., the network internally learned to propose and evaluate future candidate action sequences). While such works also show that model-free RL can produce thinking behavior, the main difference is that thinking refers to the computation done by the agent’s policy, whereas we study thinking as an action selected by the agent’s policy. More similar to this conception of thinking, the thinker architecture (Chung et al. (2023); Wang et al. (2025)) trains a policy to select actions both for environment interaction and planning in a learned model. In contrast, we consider a model-free conception of thinking.
@@
Cognition as Action
The idea of thinking as a form of action has a long history in AI and RL. Minsky theorized on the mind as a society of agents whose collective actions produce cognition and action (Minsky, 1986). Klopf’s hedonistic neuron hypothesis modelled neurons as individual RL agents (Klopf, 1982). This hypothesis has led to a line of work studying alternative neural network architectures in which artificial neural activity is produced by the stochastic policies of simple RL agents (Thomas, 2011; Kostas et al., 2019; Gupta et al., 2021). More similar to our work, some prior work has considered augmenting the agent’s environment action-space with a form of mental action. These works have focused on the challenge of memory in partially observable domains and using explicit memory read or write actions as an alternative to recurrent neural networks. For instance, Peshkin et al. (2001) augment an RL agent with a set of memory write actions and the contents of the memory are then provided to the agent as an additional input along with the environment state; Zhang et al. (2015) extend this approach to robot manipulation tasks. Various neural architectures have been developed with external memory that can be written to and read from (e.g., Graves et al., 2014; Zaremba and Sutskever, 2016) and Oh et al. (2016) explored these architectures for RL. The thought MDP model differs in that the agent has access to a Markov state representation and so thought actions serve to manipulate the agent’s policy as opposed to remembering details of the past. Prior work has used the framework of rational meta-reasoning to study computation selection as a form of action (Hay et al., 2012) and even consider using RL to learn a computation selection policy (Callaway et al., 2018). These works differ in that they study choosing among concrete computations as opposed to the abstract model of thinking as manipulating the agent’s internal state that we consider.
@@ 3.1 RL in Markov Decision Processes
RL environments are typically modeled as Markov decision processes (MDPs). Formally, an MDP is a tuple, $\langle\mathcal{S},\mathcal{A},p,r,\gamma\rangle$, where $\mathcal{S}$ is the environment state space, $\mathcal{A}$ is the set of actions that the agent can take to influence the environment state, $p:\mathcal{S}\times\mathcal{A}\rightarrow\Delta(\mathcal{S})$ is a stochastic state transition function, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a scalar-valued reward function, and $\gamma$ is a discount factor. At any moment in time, the agent is in state $S_{t}$, takes an action, $A_{t}$, receives a reward, $R_{t}=r(S_{t},A_{t})$, and transitions to a new state, $S_{t+1}$. Then the process repeats from $S_{t+1}$ until a special terminal state, $s_{\infty}$, is reached. The agent selects actions according to a policy, $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$. The value of using a policy from a particular state, $s$, is defined to be $v_{\pi}(s)\coloneqq\mathbf{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{t}|S_{0}=s]$. In RL, the agent’s objective is to find a policy that maximizes $v_{\pi}(s)$ in all states.
@@ 3.2 Thought MDPs
We now extend the MDP model to explicitly model thinking. We formally define a thought MDP as the tuple $\langle{\mathcal{S}},{\mathcal{A}},p,r,\gamma,{\mathcal{T}},{\mathcal{C}},p_{{\mathcal{T}}}\rangle$, where $\mathcal{S}$} and $\mathcal{A}$} are the environment’s state and action space respectively and $p,r,\gamma$ are defined as they are for MDPs. We add $\mathcal{T}$} as the set of thought states, $\mathcal{C}$} as the set of thought actions, and $p_{{\mathcal{T}}}:{\mathcal{S}}\times{\mathcal{T}}\times{\mathcal{C}}\rightarrow\Delta({\mathcal{T}})$ as the thought transition function. Recalling our informal definition of thinking in the introduction, we emphasize that thought states and actions do not affect environment state transitions and rewards. The agent’s objective remains to maximize cumulative discounted reward across all environment states. [^1] We base the thought MDP model on the widely used discounted problem formulation. In Appendix Appendix B, we extend the thought MDP model and theoretical results to the undiscounted, fixed-horizon setting which more closely resembles contemporary applications of RL to LLMs.
Policies in Thought MDPs
There are different ways to define the agent’s policy in a thought MDP. In this work, we formalize policies as a mapping $\pi:{\mathcal{S}}\times{\mathcal{T}}\rightarrow\Delta({\mathcal{A}}\cup{\mathcal{C}})$. This choice means that the agent can select either an environment action or a thought action at any interaction time-step but not both. We make this choice to connect to work in LLMs that inspired this model, but an alternative is that the policy is a mapping $\pi:{\mathcal{S}}\times{\mathcal{T}}\rightarrow\Delta({\mathcal{A}}\times{\mathcal{C}})$, i.e., the agent can think and act at the same time. Such modelling might be more appropriate for real-time domains where the agent cannot simply sit and think as the rest of the world evolves around it. There, of course, may be other alternatives that could be considered in future work. Below, we will use the notation $\pi(\tau)$ to refer to the agent’s state-dependent policy with the thought state fixed at $\tau$.
Interaction in Thought MDPs
In a thought MDP, interaction proceeds as follows. Episodes begin in an environment state, $S_{0}$, and thought state, $\tau_{0}$, and the agent chooses either an environment action or thought action according to its policy. Let $A_{0}$ be the random variable denoting this choice. If $A_{0}\in{\mathcal{A}}$ then the environment state, $S_{1}$, at the next time-step is sampled from $p(\cdot|S_{0},A_{0})$, the thought state $\tau_{1}$ keeps the value of $\tau_{0}$, and the agent receives reward $R_{0}\coloneqq r(S_{0},A_{0})$. Conversely, if $A_{0}\in{\mathcal{C}}$, then $S_{1}$ keeps the value of $S_{0}$, $\tau_{t}$ is sampled from $p_{{\mathcal{T}}}(\cdot|S_{0},\tau_{0},C_{0})$, and the agent receives $R_{0}=0$. Then the process repeats with the agent selecting either an environment or thought action from its policy until the agent reaches a special terminal state $s_{\infty}\in{\mathcal{S}}$. For thought MDPs, we define the value function to be $v_{\pi}(s,\tau)\coloneqq\mathbf{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{t}|S_{0}=s,\tau_{0}=\tau]$.
Key Assumptions
@@
Our formalism raises questions about how time should be treated for the two forms of action. First, should the discount factor be applied equally for both thinking and non-thinking time-steps? Equal application discourages thinking as thought actions do not influence reward either directly or indirectly through the environment state. Nevertheless, as we shall see, thought actions still might be selected if they ultimately cause the agent to choose a better environment action. Alternatively, we could apply a different discount factor for thinking time-steps to reflect the actual time-delay of thinking compared to acting in a given domain. For example, if thinking takes $(1/k)$ the time of any environment action then we could use a discount of $\gamma^{\frac{1}{k}}$. In this work, we will assume that the discount is applied the same at both thinking and non-thinking time steps. The second related issue is that the proposed model assumes that the environment state remains constant while the agent takes thought actions. Such an assumption might be reasonable for relatively static environments (such as generating text or board games) but is problematic in dynamic environments (such as autonomous driving) where waiting to act can be consequential. Again, for simplicity, we will focus on the static environment case, but note this issue as another interesting direction to refine the model.
@@
[…]
@@ 4 Policy Initialization Determines Emergence of Thinking
For exposition’s sake, suppose ${\mathcal{T}}=\{\tau_{0},\tau_{1}\}$, in which case we can view the agent’s policy, $\pi$, as consisting of two sub-policies, $\pi(\tau_{0})$ and $\pi(\tau_{1})$. Now consider the example depicted in Figure˜1 where $\pi(\tau_{0})$ is initialized sub-optimally (always takes an action that leads away from the goal) while $\pi(\tau_{1})$ is initialized to be optimal. We show the progress of policy iteration in this domain in Figure˜1. After the first iteration of policy improvement, the policy has learned to first change the agent’s thought state to $\tau_{1}$ (top row) from which it then follows $\pi(\tau_{1})$ to reach the goal. After four iterations, in environment states close to the goal, the policy just directly moves right without changing its thought state while continuing to first take a thought action in states that are far from the goal. Finally, after ten iterations, the policy converges to the optimal policy and simply moves to the right without thinking. This example shows that, while thinking is sub-optimal in the long-run, it can be beneficial early in learning by allowing the agent to use sub-policies already contained in its policy function.
@@
Policy iteration sets $\pi^{\prime}(s,\tau)\leftarrow\arg\max_{a\in{\mathcal{A}}\cup{\mathcal{C}}}q_{\pi}(s,\tau,a)$ where
$q_{\pi}(s,\tau,a)\coloneqq\begin{cases}r(s,a)+\gamma\mathbf{E}_{{s}^{\prime}}[v_{\pi}({s}^{\prime},\tau)]&\text{if }a\in{\mathcal{A}}\\ \gamma v_{\pi}(s,{\tau}^{\prime})&\text{if }a\in{\mathcal{C}}.\end{cases}$
@@
Thus, a thought action is selected when the policy is such that changing the thought state to $\tau$$^{\prime}$ will lead to a greater expected return from $s$ than when leaving the thought state at $\tau$. ∎
Although, Theorem˜5 does not directly say anything about policy initialization, the condition for a thought action to be selected requires $\pi(\tau)$ and $\pi({\tau}^{\prime})$ to be somehow set up as different such that there could be an advantage in shifting from $\tau$ to $\tau$$^{\prime}$. Thus, policy initialization is critical for whether thinking will emerge or not.
Theorem˜5 deals with exact policy iteration and does not address the impact of policy initialization on the emergence of thinking when learning from samples. We next turn to the concept of the effective horizon introduced by Laidlaw et al. (2023). The effective horizon was introduced as a problem complexity parameter for MDPs that aligns with the observed difficulty of common RL benchmarks. Laidlaw et al. then derived sample complexity bounds for model-free RL algorithms in terms of the effective horizon and showed that these bounds were predictive of the success of deep RL algorithms on benchmark problems. We will build upon their specific result on goal MDPs.
@@
A goal (thought) MDP is an (thought) MDP with a set of absorbing environment states, $\mathcal{S}_{\mathtt{goal}}\subset{\mathcal{S}}$, and $r(s,a)=1$ if $s\in\mathcal{S}_{\mathtt{goal}}$ and 0 otherwise.
For goal MDPs, Laidlaw et al. (2023) showed that the effective horizon, $H$ can be upper-bounded by $1+\log_{\mathcal{A}}\frac{\log 2l}{p}$ where $l$ is a maximum episode length, $p\leq\Pr_{\mathtt{expl}}(s_{T}\in\mathcal{S}_{\mathtt{goal}}|s_{t},a_{t})$ is a lower bound on the probability of the initial exploration policy, $\pi_{\mathtt{expl}}$, discovering a goal state after taking action $a$ in state $s$ at any time-step. Intuitively, if $p$ is larger, then an RL algorithm has an easier time discovering rewarding action sequences. We next show that thought actions can be used to reduce the effective horizon of a given instance. Note that Laidlaw et al. (2023) proved their results assuming deterministic MDPs, so we also adopt this assumption for the next proposition.
@@
Suppose we have a goal thought MDP that has ${\mathcal{T}}=\{\tau_{0},\tau_{1}\}$. Let the initial policy, $\pi$, be such that there is some $p_{0}$ such that $\Pr(s_{l}\in s_{\mathtt{goal}}|s,a,\pi(\tau_{0}))>p_{0}$ for all $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$ and some $p_{1}$ such that $\Pr(s_{l}\in s_{\mathtt{goal}}|s,a,\pi(\tau_{1}))>p_{1}$ for all $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$. Let $c_{s}$ be the thought action such that $\tau_{1}=p_{{\mathcal{T}}}(s,\tau_{0},c_{s})$ and let $p_{c}$ lower bound the probability $\pi(c_{s}|s,\tau_{0})$. If $p_{c}\cdot p_{1}>p_{0}$, then thought actions reduce the upper bound on the effective horizon of the MDP.
See Appendix˜A for the proof. Intuitively, policy $\pi(\tau_{1})$ has a higher probability of finding a goal state from any state compared to simply running $\pi(\tau_{0})$. If the probability of taking a thought action that changes $\tau_{0}$ to $\tau_{1}$ is not too low then it is easier to find a goal state by first changing the thought state and then executing $\pi(\tau_{1})$. Again, this result suggests that the benefit of thinking depends crucially on policy initialization.
@@
One interpretation of Theorem˜5 is that thought actions can function as policy improvement operators applied to a particular state. This interpretation is interesting as it aligns with the use of decision-time planning in RL to refine an agent’s choice of action in a way that focuses computation on its current state (Sutton and Barto, 2018). While thought actions do not involve look-ahead search with a model, Theorem˜5 shows that their utility is also in providing local policy improvement. Theorem˜5 shows this utility for the choice of a single thought action and we also present a corollary showing that if policy improvement produces a policy that thinks for consecutive steps in $s$ then each thinking step will further improve upon the action $\pi(s,\tau)$.